3. Given an Epicyclic gearbox that has a fixed outer gear “C”, with 400 Teeth, Planet gears have 40 Teeth, the input is Arm “ please show all work
Answer #1

Solution Given to = 400 to = 40 C400) 46) D020) arm fig: epicyclic planet fsun gear train arm Sun 1 pt planet - to 4 to gedyAs rc = 2 Rot Ro to 400 2 =40+ 2 to 320 m = 1200 rpm C Assumed counter clockwise tue to no to 1200 - 320 no 400 n = 1200X400Ti → Input torque T2 → output torgue Io → Holding torquel fixing torque Tit Tzt T3 =0 pout = n Pin :: T2W2= -0.8 X T3 W3 .: T